# Search for Ultralight Dark Matter with Quantum Magnetometry in the Earth’s Cavity

### v1

@@ Abstract
Ultralight dark matter candidates, such as axions and dark photons, are leading dark matter candidates. They may couple feebly to photons, sourcing oscillating electromagnetic signals in the presence of magnetic field. The Earth resonant cavity formed between the ground and the ionosphere provides a natural waveguide that can amplify such signals. We carry out a project aiming to search for new physics using the unshielded high-sensitivity atomic magnetometer, termed the Geomagnetic Probe for nEw physiCS (GPEX). In this work, we report our first search for axion and dark photon dark matter, conducted in the desert of XiaoDushan in Gansu Province, China. Analysis of the collection of one-hour data shows no evidence for axion- or dark photon-induced magnetic signals. Correspondingly, we set the constraints on the axion-photon coupling with $g_{a\gamma\gamma}<7\times 10^{-10}\,\mathrm{GeV^{-1}}$ and the dark photon kinetic-mixing parameter $\epsilon<2\times 10^{-6}$ in the mass range $3.5\times 10^{-16}\,\mathrm{eV}\sim 1.8\times 10^{-14}\,\mathrm{eV}$. Our findings demonstrate the feasibility of using ground-based quantum magnetic sensors for ultralight dark matter searches. Future networks of such detectors operating over extended periods could further enhance sensitivity, approaching the limits set by astrophysical observations.
Introduction. Cosmological observations indicate that only about 16% of the universe’s total matter is baryonic. While the Standard Model of particle physics accurately describes this component, the nature of the remaining $\sim$84%—the enigmatic dark matter—remains a fundamental problem at the intersection of particle physics, cosmology, and astrophysics Silk et al. (2010). Consequently, leading dark matter candidates now arise from extensions to the Standard Model that introduce ultralight fields, such as axions, axion-like particles, and dark photons Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Nelson and Scholtz (2011); Arias et al. (2012); Jaeckel (2012).
@@
The axion, originally proposed to solve the strong-CP problem in quantum chromodynamics Peccei and Quinn (1977), also emerges naturally in high-energy frameworks such as string theory and extra-dimensional models Arvanitaki et al. (2010); Svrcek and Witten (2006); Abel et al. (2008a). Another promising candidate, the dark photon, provides a portal to the dark sector through kinetic mixing with the Standard Model photon Holdom (1986); Dienes et al. (1997); Abel et al. (2008b); Goodsell et al. (2009); Nelson and Scholtz (2011). A key cosmological virtue of these ultralight fields is that they can be produced in the early universe via non-thermal mechanisms, naturally accounting for the observed dark matter relic density Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016).
@@
Ultralight dark matter (ULDM), with masses ranging from $10^{-22}\,\mathrm{eV}$ to a few $\mathrm{eV}$, behaves as a weakly coupled, coherent classical field that permeates all space. Probing ultralight axions and dark photons relies on their non-gravitational couplings to Standard Model particles Bradley et al. (2003); Budker et al. (2014); An et al. (2013); Anastassopoulos et al. (2017); Braine et al. (2020); Gramolin et al. (2021); Smorra et al. (2019); Jiang et al. (2021); Chen et al. (2022a); Aprile et al. (2022); Bloch et al. (2023); Wei et al. (2025); Jiang et al. (2024); Arza et al. (2024); An et al. (2024); Gavilan-Martin et al. (2025). Consequently, a suite of specialized experimental approaches has been developed to probe their couplings to the photon field. These include shielded experiments such as ADMX Braine et al. (2020), DM-Radio Brouwer et al. (2022), ABRACADABRA Salemi et al. (2021), and others An et al. (2024); Kimball and van Bibber (2023); Sikivie (2021); Irastorza and Redondo (2018). On the other hand, experiments involving quantum sensing and their networks, such as atomic spectroscopy Berger and Bhoonah (2022), superconducting cavities Dixit et al. (2021); Agrawal et al. (2024); Cervantes et al. (2024); Tang et al. (2024); Zeng et al. (2025); Zheng et al. (2025), atomic magnetometers Afach et al. (2021); Jiang et al. (2024), state-swapping techniques Chen et al. (2022b); Wurtz et al. (2021); Chen et al. (2025); Jiang et al. (2023), and squeezed-state receivers Backes et al. (2021), have achieved unprecedented sensitivities. These advances are attributed to their ability to enhance the effective sensitivity of the sensors and to confidently distinguish signals from various sources of noise using correlated data.
@@
In this work, we utilize an unshielded atomic magnetometer—an optically pumped rubidium magnetometer—to conduct the Geomagnetic Probe for EXotic physics (GPEX), targeting dark matter axions and dark photons. The experiment was carried out in the desert of XiaoDushan, Gansu Province, China, a location far from anthropogenic electromagnetic noise. By leveraging the Earth as a giant transducer, the natural conducting cavity formed by the Earth’s surface and ionospheric layers facilitates the conversion of ultralight dark matter fields into detectable magnetic field signals. This enhancement is particularly effective for ultralight axion and dark photon fields with Compton wavelengths comparable to or larger than the Earth’s size, corresponding to a mass range of $3.5\times 10^{-16}\,\mathrm{eV}\sim 1.8\times 10^{-14}\,\mathrm{eV}$, where the Earth’s radius $R_{e}\gg L$ (with $L$ being the characteristic size of the field oscillation). We analyze one hour of data collected using our high-sensitivity quantum magnetometer and identify no robust signal candidates from dark matter axions and dark photons. Based on this null result, we derive constraints on the axion-photon coupling $g_{a\gamma\gamma}$ and the dark photon kinetic mixing parameter $\epsilon$ within the specified mass range. Owing to the exceptional sensitivity of the magnetometer and the quiet experimental environment, our limits reach $g_{a\gamma\gamma}\sim 7\times 10^{-10}\,\mathrm{GeV^{-1}}$ and $\epsilon\sim 2\times 10^{-6}$. These results demonstrates the strong potential of unshielded atomic magnetometers for rare event searches.
Theoretical background. Ultralight dark matter can interact feebly with electromagnetic fields. Because its occupancy number is extremely high, the dark matter field behaves classically. This collective effect can be incorporated into Maxwell’s equations as an effective classical current $\vec{J}_{\mathrm{DM}}$, which acts as a source for electromagnetic phenomena. In this context, the phenomenology is reduced to solve the classical equations Sikivie (2021)
@@
where $\ell=0,1,2,$... and $m=-\ell,-\ell+1,...,\ell-1,\ell$. The parameter $\omega$ corresponds to the typical angular frequency of the system, which is equivalent to the dark matter particle masses $m_{a}$ and $m_{\gamma^{\prime}}$ for axions and dark photons respectively. For ultralight fields such as axions and dark photons, the identity $\vec{\nabla}\times\vec{J}_{\mathrm{DM}}=0$ is satisfied. From VSH properties, it is implied directly that $J_{\ell m}^{(2)}=0$, and that $J_{\ell m}^{(r)}=r{J_{\ell m}^{(1)\prime}}+J_{\ell m}^{(1)}$. We look for magnetic field solutions at the Earth’s surface, considering that electromagnetic fields are bounded on the one side by the ground and on the other side by the ionosphere. For Compton wavelengths bigger than the ionosphere height, the dark matter induced magnetic field signal $\vec{B}_{\mathrm{DM}}$ at the Earth’s surface is determined uniquely by the radial component $J_{\ell m}^{(r)}$ of the effective current. We find
@@
For axions, the effective current at $r=R_{e}$ is Arza et al. (2022)
@@
where $\epsilon$ is the kinetic mixing parameter between dark photons and standard model photons and $A_{m}^{\prime}$ the dark photon amplitudes defined in terms of $\vec{A}^{\prime}=(A_{x}^{\prime},A_{y}^{\prime},A_{z}^{\prime})$ as $A_{1}=-{1\over\sqrt{2}}(A_{x}^{\prime}-iA_{y}^{\prime})$, $A_{1}={1\over\sqrt{2}}(A_{x}^{\prime}+iA_{y}^{\prime})$ and $A_{0}^{\prime}=A_{z}^{\prime}$, where $\vec{A}^{\prime}$ is normalized as ${1\over 2}m_{\gamma^{\prime}}^{2}\left<|\vec{A}|^{2}\right>=\rho_{\mathrm{0}}$. For dark photons, the oscillation frequency is defined as $\omega_{m}=m_{\gamma^{\prime}}-2\pi mf_{d}$, where $f_{d}$ is the frequency associated to the sidereal day, which accounts for the effects of Earth’s rotation. As the data used in this work correspond to a couple of hours, we neglect Earth’s rotation effects. Just like in the axion case, we identify $J_{\ell m}^{(r)}$ in Eq. (9) and find a dark photon magnetic field signal given by
@@
We analyze the data by searching for excess power in the frequency domain, as characterized by the power spectral density (PSD) Groth (1975); Derevianko (2018); Fedderke et al. (2021b). The remote desert location of the experiment minimizes dominant anthropogenic noise contributions at low frequencies. The primary source of noise in the recorded data is the magnetometer itself, which we assume to exhibit both Gaussian and stationary behavior throughout the measurement period. Below approximately 0.5 Hz, the noise follows a $1/f$ power-law dependence, while at higher frequencies, it becomes approximately frequency-independent. Accordingly, we restrict our analysis to the frequency range between 0.1 Hz and 5 Hz. This range is chosen to ensure reliable extraction of magnetic field signals under the condition $(\omega R_{e})^{2}<2$, where $\omega$ is the angular frequency and $R_{e}$ is the Earth’s radius. Within this band, the noise variance is estimated from the power in adjacent frequency bins at each frequency. This approach yields a nearly flat noise background, providing an optimal regime in which narrow spectral features—such as potential signals from ultralight dark matter—can be identified and studied.
@@
where we have approximated $\delta(0)\approx T_{\rm obs}$ for finite observation time. The noise PSD, $S_{\rm n}(f_{\rm s})$, follows a $\chi^{2}$ distribution with two degrees of freedom Groth (1975). For cases where the observation time exceeds the coherence time, we refer the reader to Refs. Fedderke et al. (2021b); Nishizawa et al. (2025) for the definition of the weighted estimator applied over multiple data segments, which incorporates the signal bandwidth and corresponding corrections.
Figure 4: Left: The 95% confidence level (CL) exclusion limit on the axion-photon coupling $g_{a\gamma\gamma}$ as a function of the axion mass $m_{a}$, derived from the first run of the GPEX experiment. We also display the relevant experimental bounds based on the SuperMAG analysis Arza et al. (2022); Friel et al. (2024), the SNIPE-hunt analysis Sulai et al. (2023), the Eskdalemuir observation Taruya et al. (2025a); Nishizawa et al. (2025), and the CAST result Anastassopoulos et al. (2017). The prospective sensitivity, evaluated using a network of 10 magnetometers, each with a sensitivity of 0.54 $\rm{pT/\sqrt{Hz}}$ and operating for one month, is also shown. Right: The 95% CL exclusion limit on the dark photon kinetic-mixing parameter $\epsilon$ as a function of the dark photon mass $m_{\gamma^{\prime}}$, derived from the first run of the GPEX experiment. We show the limits from the SuperMAG Fedderke et al. (2021a); Friel et al. (2024) and SNIPE-hunt Sulai et al. (2023) analyses based on geomagnetic field observations, as well as the limit from the AMAILS experiment Jiang et al. (2024) based on a synchronized quantum sensor network. The prospective sensitivity is evaluated under the same configuration as for the axion case (i.e., a network of 10 magnetometers, each with 0.54 $\rm{pT/\sqrt{Hz}}$ sensitivity, operating for one month).
@@
Our results are presented in Fig. 4, which shows the 95% CL exclusion limits on the couplings $g_{a\gamma\gamma}$ and $\epsilon$ over the frequency range $0.1\leq f\leq 5\,\mathrm{Hz}$. The corresponding mass range for axion or dark photon dark matter is from $3.5\times 10^{-16}\,\mathrm{eV}$ to $1.8\times 10^{-14}\,\mathrm{eV}\;$. In the same figure, we also compare our limits with existing constraints from geomagnetic field observations conducted by the SuperMAG Arza et al. (2022); Fedderke et al. (2021a, b); Arza et al. (2025), SNIPE-hunt Sulai et al. (2023), and Eskdalemuir Observatory Nishizawa et al. (2025) collaborations, which employed various types of classical magnetometers.
For the axion, our constraint is approximately one order of magnitude stronger than that of the SNIPE-hunt experiment, primarily due to the high sensitivity of our quantum magnetometer. However, our limits are weaker than those from SuperMAG and Eskdalemuir, mainly owing to their significantly longer observation times. The results from the CAST solar axion search Anastassopoulos et al. (2017) are also included in the figure. We note that additional astrophysical observations Yuan et al. (2021); Caputo et al. (2022); Wouters and Brun (2013); Dessert et al. (2020); Marsh et al. (2017), including X-ray observations of the quasar H1821+643 obtained with the Chandra telescope Reynés et al. (2021), provide even stronger constraints on the axion within the same mass and frequency range.
For the dark photon, our constraint is more sensitive than those from SNIPE-hunt and AMAILS Jiang et al. (2024), which utilized a synchronized quantum sensor network. However, our limits are not as strong as those from SuperMAG, primarily due to the shorter duration of our observation and the smaller number of magnetometers deployed. Astrophysical constraints Fischbach et al. (1994); Marocco (2021); Bhoonah et al. (2019), particularly those derived from dark photon-induced heating and cooling in the gas-rich Leo T dwarf galaxy Wadekar and Farrar (2021), yield the most stringent limits. Nevertheless, these astrophysical bounds rely strongly on the inferred cooling rates of the relevant systems and the detailed understanding of the underlying cooling mechanisms.
@@
We aim to surpass the SuperMAG and Eskdalemuir limits derived from long-term monitoring by conducting extended observations with more sensitive quantum magnetometers. As the current experimental reach is constrained by magnetometer sensitivity rather than geomagnetic noise, we plan to establish a high-performance network of quantum magnetometers to search for dark matter signals and potentially exceed existing astrophysical bounds. Such a network can enhance the effective sensitivity and, through correlated measurements, enable confident discrimination of dark matter signals from various noise sources Pospelov et al. (2013); Derevianko and Pospelov (2014); Roberts et al. (2017); Afach et al. (2021); Jiang et al. (2024). For instance, employing 10 quantum magnetometers with sensitivity of 0.54 $\rm{pT/\sqrt{Hz}}$ Huang et al. (2025), improved by a factor of 40, could enhance the overall sensitivity by about three orders of magnitude after one month of integration, as shown in Fig. 4. In the higher-frequency regime, signal model predictions are available Taruya et al. (2025b), allowing our search to be extended beyond the low-frequency range. Furthermore, as mentioned in Sulai et al. (2023), employing local multi-sensor arrays to measure the curl of the local magnetic field can extend the accessible frequency range to approximately 1 kHz.
Acknowledgement. This work is supported by the National Key Research and Development Program of China (Grant No. 2022YFA1605000), and Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2024ZD0302300). L.W. is supported by the NNSFC under Grant No. 12275134 and No. 12335005. J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China (NNSFC) under Grants No. 12025507, No.12450006. J.G. is supported by the NNSFC under Grant No. 12305111. B.Z. is supported by the NNSFC under Grant No. 12275232 and Shangdong Provincial Natural Science Foundation for Distinguished Young Scholars under Grant No. ZR2025QA20. Q.Y. is supported by the Project for Young Scientists in Basic Research of Chinese Academy of Sciences under Grant No. YSBR-061.
@@
Ariel Arza,1,2 Yuanlin Gong,1 Jun Guo,3 Xiaofei Huang,4 Jing Shu,5,6,7 Hongliang Tian,1 Wenyu Wang,8 Kai Wei,4,9 Lei Wu,1,2 Mingming Xia,10 Jin Min Yang,11,12 Qiang Yuan,13,14 Yang Zhang11 Yi Zhang,13,14 Bin Zhu,2
@@
13Key Laboratory of Dark Matter and Space Astronomy, Purple Mountain Observatory, Chinese Academy of Sciences, Nanjing 210023, China

### v2

@@ Abstract
Ultralight dark matter candidates, such as axions and dark photons, are leading dark matter candidates. They may couple feebly to photons, sourcing oscillating electromagnetic signals in the Earth’s conducting cavity formed between the ground and the ionosphere, providing detectable magnetic field signatures at wavelengths above the Earth’s size. We carry out a project aiming to search for new physics using an unshielded high-sensitivity atomic magnetometer, termed the Geomagnetic Probe for nEw physiCS (GPEX). In this work, we report our first search for axion and dark photon dark matter, conducted in the desert of XiaoDushan in Gansu Province, China. Analysis of the collection of one-hour data shows no robust evidence for axion- or dark photon-induced magnetic signals. Correspondingly, we set the constraints on the axion-photon coupling with $g_{a\gamma\gamma}<7\times 10^{-10}\,\mathrm{GeV^{-1}}$ and the dark photon kinetic-mixing parameter $\epsilon<2\times 10^{-6}$ in the mass range $3.5\times 10^{-16}\,\mathrm{eV}\sim 1.8\times 10^{-14}\,\mathrm{eV}$. Our findings demonstrate the feasibility of using ground-based quantum magnetic sensors for ultralight dark matter searches. Future networks of such detectors operating over extended periods could improve the sensitivity by about three orders of magnitude.
@@
Introduction. Cosmological observations indicate that only about 16% of the universe’s total matter is baryonic. While the Standard Model of particle physics accurately describes this component, the nature of the remaining $\sim$84%—the enigmatic dark matter—remains a fundamental problem at the intersection of particle physics, cosmology, and astrophysics Silk et al. (2010). Consequently, ultralight dark matter (ULDM) candidates now arise from extensions to the Standard Model, such as axions, axion-like particles, and dark photons Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Nelson and Scholtz (2011); Arias et al. (2012); Jaeckel (2012). The axion, originally proposed to solve the strong-CP problem in quantum chromodynamics Peccei and Quinn (1977), also emerges naturally in high-energy frameworks such as string theory and extra-dimensional models Arvanitaki et al. (2010); Svrcek and Witten (2006); Abel et al. (2008a). Another promising candidate, the dark photon, provides a portal to the dark sector through kinetic mixing with the Standard Model photon Holdom (1986); Dienes et al. (1997); Abel et al. (2008b); Goodsell et al. (2009); Nelson and Scholtz (2011). A key cosmological virtue of these ultralight fields is that they can be produced in the early universe via non-thermal mechanisms, naturally accounting for the observed dark matter relic density Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016).
@@
ULDM with masses ranging from $10^{-22}\,\mathrm{eV}$ to a few $\mathrm{eV}$, behaves as a weakly coupled, coherent classical field that permeates all space. Detection of ULDM relies on their non-gravitational couplings to Standard Model particles Bradley et al. (2003); Budker et al. (2014); An et al. (2013); Anastassopoulos et al. (2017); Irastorza and Redondo (2018); Braine et al. (2020); Gramolin et al. (2021); Smorra et al. (2019); Sikivie (2021); Jiang et al. (2021); Chen et al. (2022a); Aprile et al. (2022); Bloch et al. (2023); Wei et al. (2025); Jiang et al. (2024); Arza et al. (2024); An et al. (2024); Kimball and van Bibber (2023); Xu et al. (2024); Gavilan-Martin et al. (2025). For instance, a suite of specialized experimental approaches has been developed to probe their couplings to the photon field. These include shielded experiments such as ADMX Braine et al. (2020), DM-Radio Brouwer et al. (2022), ABRACADABRA Salemi et al. (2021), and others involving quantum sensing and their networks, such as atomic spectroscopy Berger and Bhoonah (2022), superconducting cavities Dixit et al. (2021); Agrawal et al. (2024); Cervantes et al. (2024); Tang et al. (2024); Zeng et al. (2025); Zheng et al. (2025), atomic magnetometers Afach et al. (2021); Jiang et al. (2024), state-swapping techniques Chen et al. (2022b); Wurtz et al. (2021); Chen et al. (2025); Jiang et al. (2023), and squeezed-state receivers Backes et al. (2021). Alternatively, the Earth’s interior and ionosphere forms a natural cavity that functions as a giant transducer for ULDM. In this cavity, the conversion of dark matter into a detectable magnetic field is highly effective for signal frequencies below 7 Hz, which corresponds to dark matter masses less than $3\times 10^{-14}\,\mathrm{eV}$ (Compton wavelengths larger than the Earth Radius). Competitive constraints on axion and dark photon dark matter have been derived from the analysis of long-term geomagnetic datasets from the SuperMAG collaboration Arza et al. (2022); Fedderke et al. (2021a, b); Friel et al. (2024) and the Eskdalemuir Observatory (UK) Nishizawa et al. (2025); Nomura et al. (2025). However, conventional long-term geomagnetic measurements usually suffer from systematics and are not optimized for dark matter searches. On the other hand, the SNIPE-Hunt collaboration has further advanced the search by setting constraints with a dedicated network of classical magnetometers Sulai et al. (2023); hig . Building on this idea, future efforts can extend to different search strategies Bloch and Kalia (2024) and the probing of other ULDM candidates such as millicharged particles Arza et al. (2025).
@@
This work reports the inaugural results of the Geomagnetic Probe for nEw physiCS (GPEX) experiment, which searches for geomagnetic signatures from ULDM candidates. For this first-stage campaign, we operated an unshielded, optically pumped rubidium vapor magnetometer in the electromagnetically pristine environment of the XiaoDushan desert in China. The choice of quantum sensor enables precise measurements and can be readily integrated into future networks of synchronized sensors, while the remote desert environment minimizes anthropogenic electromagnetic disturbances, enhancing our ability to detect subtle signals. Our analysis of one hour of data revealed no robust signals of axions or dark photons. We therefore place upper limits on their couplings to photons, achieving constraints of $g_{a\gamma\gamma}\sim 7\times 10^{-10}\,\mathrm{GeV^{-1}}$ for axions and $\epsilon\sim 2\times 10^{-6}$ for dark photons, which surpass those from the previous SNIPE-Hunt collaboration by an order of magnitude and a factor of three, respectively. This demonstrates the powerful potential of unshielded atomic magnetometers for this search.
Theoretical background. Ultralight dark matter can interact feebly with electromagnetic fields. Because its occupancy number is extremely high, the dark matter field behaves classically. This collective effect can be incorporated into Maxwell’s equations as an effective classical current density $\vec{J}_{\mathrm{DM}}$, which acts as a source for electromagnetic phenomena. In this context, the phenomenology is reduced to solve the classical equations Sikivie (2021)
@@
where $\ell=0,1,2,$... and $m=-\ell,-\ell+1,...,\ell-1,\ell$. The parameter $\omega$ corresponds to the typical angular frequency of the system, which is equivalent to the dark matter particle masses $m_{a}$ and $m_{\gamma^{\prime}}$ for axions and dark photons, respectively. For ultralight fields such as axions and dark photons, the identity $\vec{\nabla}\times\vec{J}_{\mathrm{DM}}=0$ is satisfied. From VSH properties, it is implied directly that $J_{\ell m}^{(2)}=0$, and that $J_{\ell m}^{(r)}=r{J_{\ell m}^{(1)\prime}}+J_{\ell m}^{(1)}$. We look for magnetic field solutions at the Earth’s surface, considering that electromagnetic fields are bounded on the one side by the ground and on the other side by the ionosphere. For Compton wavelengths bigger than the ionosphere height, the dark matter induced magnetic field signal $\vec{B}_{\mathrm{DM}}$ at the Earth’s surface is determined uniquely by the radial component $J_{\ell m}^{(r)}$ of the effective current. We find
@@
For axions, the effective current at $r=R_{e}$ is Arza et al. (2022); Taruya et al. (2025a)
@@
where $\epsilon$ is the kinetic mixing parameter between dark photons and standard model photons and $A_{m}^{\prime}$ the dark photon amplitudes defined in terms of $\vec{A}^{\prime}=(A_{x}^{\prime},A_{y}^{\prime},A_{z}^{\prime})$ as $A_{1}=-{1\over\sqrt{2}}(A_{x}^{\prime}-iA_{y}^{\prime})$, $A_{-1}={1\over\sqrt{2}}(A_{x}^{\prime}+iA_{y}^{\prime})$ and $A_{0}^{\prime}=A_{z}^{\prime}$, where $\vec{A}^{\prime}$ is normalized as ${1\over 2}m_{\gamma^{\prime}}^{2}\left<|\vec{A}|^{2}\right>=\rho_{\mathrm{0}}$. For dark photons, the oscillation frequency is defined as $\omega_{m}=m_{\gamma^{\prime}}-2\pi mf_{d}$, where $f_{d}$ is the frequency associated to the sidereal day, which accounts for the effects of Earth’s rotation. As the data used in this work correspond to a couple of hours, we neglect Earth’s rotation effects. Just like in the axion case, we identify $J_{\ell m}^{(r)}$ in Eq. (9) and find that the dark photon magnetic field signal is given by
@@
We analyze the data by searching for excess power in the frequency domain, as characterized by the power spectral density (PSD) Groth (1975); Derevianko (2018); Fedderke et al. (2021b). The remote desert location of the experiment minimizes dominant anthropogenic noise contributions at low frequencies. The primary source of noise in the recorded data is the magnetometer itself, which we assume to exhibit both Gaussian and stationary behavior throughout the measurement period. Below approximately 0.5 Hz, the noise follows a $1/f$ power-law dependence, while at higher frequencies, it becomes approximately frequency-independent. Accordingly, we restrict our analysis to the frequency range between 0.1 Hz and 5 Hz. This range is chosen to ensure reliable extraction of magnetic field signals under the condition $(\omega R_{e})^{2}<2$, where $\omega$ is the angular frequency and $R_{e}$ is the Earth’s radius. Within this band, the noise variance is estimated from the power in adjacent frequency bins at each frequency. This yields a nearly flat noise background, providing an optimal regime in which narrow spectral features—such as potential signals from ultralight dark matter—can be identified and studied.
@@
where we have approximated $\delta(0)\approx T_{\rm obs}$ for finite observation time. The noise PSD, $S_{\rm n}(f_{\rm s})$, follows a $\chi^{2}$ distribution with two degrees of freedom Groth (1975). For cases where the observation time exceeds the coherence time, we refer the reader to Refs. Fedderke et al. (2021b); Nishizawa et al. (2025) for the definition of the weighted estimator applied over multiple data segments, which incorporates the signal bandwidth and corresponding corrections. Moreover, we provide representative examples of the noise amplitude distributions at selected frequencies in the Supplementary Material, demonstrating that they are well described by a $\chi^{2}$ distribution as expected. The $p$-values from the $\chi^{2}$ goodness-of-fit test across the entire frequency range of interest are also shown.
@@
Figure 4: Left: 95% CL limit on $g_{a\gamma\gamma}$ vs $m_{a}$ from GPEX, compared with SuperMAG Arza et al. (2022); Friel et al. (2024), SNIPE-Hunt Sulai et al. (2023), Eskdalemuir Taruya et al. (2025b); Nishizawa et al. (2025), CAST Anastassopoulos et al. (2017), and the quasar H1821+643 bound Reynés et al. (2021). Right: 95% CL limit on $\epsilon$ vs $m_{\gamma^{\prime}}$ from GPEX, compared with SuperMAG Fedderke et al. (2021a, b); Friel et al. (2024), SNIPE-Hunt Sulai et al. (2023), AMAILS Jiang et al. (2024), and the Leo T dwarf bound Wadekar and Farrar (2021). Prospective sensitivity for 10 magnetometers ($0.54\penalty 10000\ \mathrm{pT/\sqrt{Hz}}$, 1 month) is also shown.
Our results, presented in Fig. 4, shows the 95% CL exclusion limits on the couplings $g_{a\gamma\gamma}$ and $\epsilon$ over the frequency range $0.1\leq f\leq 5\,\mathrm{Hz}$. This corresponds to an axion or dark photon mass range from $3.5\times 10^{-16}\,\mathrm{eV}$ to $1.8\times 10^{-14}\,\mathrm{eV}\;$. Compared to existing constraints from geomagnetic field observations, SuperMAG Arza et al. (2022); Fedderke et al. (2021a, b), SNIPE-Hunt Sulai et al. (2023), Eskdalemuir Observatory Nishizawa et al. (2025), our bounds on axions are approximately one order of magnitude stronger than those from SNIPE-Hunt. However, they are less stringent than limits set by CAST Anastassopoulos et al. (2017), SuperMAG, and Eskdalemuir, primarily due to their substantially longer observation times. For dark photons, our constraints improve upon those from SNIPE-Hunt and AMAILS Jiang et al. (2024), yet remain weaker than the SuperMAG Fedderke et al. (2021a); Friel et al. (2024) results for similar reasons. We note that concurrent with the completion of this manuscript, new experimental constraints on dark photons have been reported by two independent groups: one analyzing archival data from the Eskdalemuir observatory Nomura et al. (2025), and another utilizing a dedicated laboratory experiment with a two-sensor array of optically pumped scalar magnetometers Zhao et al. (2025).
Additional astrophysical observations provide even more stringent constraints for both axions and dark photons in this parameter space Yuan et al. (2021); Caputo et al. (2022); Wouters and Brun (2013); Dessert et al. (2020); Marsh et al. (2017); Guo et al. (2024); Song et al. (2025); Xue et al. (2024); Guo et al. (2025); Fischbach et al. (1994); Marocco (2021); Bhoonah et al. (2019); Su et al. (2022). For axions, limits from X-ray observations of the quasar H1821+643 with the Chandra telescope are particularly strong Reynés et al. (2021). For dark photons, the most stringent constraints currently come from analyses of heating and cooling effects induced in the gas-rich Leo T dwarf galaxy Wadekar and Farrar (2021). It is important to note, however, that these astrophysical bounds depend significantly on the modeling of the relevant systems and underlying mechanisms, and thus entail substantial uncertainties.
@@
As the current experimental reach is constrained by magnetometer sensitivity rather than geomagnetic noise, we plan to establish a high-performance network of quantum magnetometers to search for dark matter signals and achieve world leading sensitivity. Such a network can enhance the effective sensitivity and, through correlated measurements, enable confident discrimination of dark matter signals from various noise sources Pospelov et al. (2013); Derevianko and Pospelov (2014); Roberts et al. (2017); Afach et al. (2021); Jiang et al. (2024). For instance, employing 10 quantum magnetometers with sensitivity of 0.54 $\rm{pT/\sqrt{Hz}}$ Huang et al. (2025), improved by a factor of 40, could enhance the overall sensitivity by about three orders of magnitude after one month of integration, as shown in Fig. 4. We anticipate that a future network of quantum magnetometers in remote locations will have the sensitivity to surpass constraints set by SuperMAG, Eskdalemuir, and eventually even astrophysical bounds. Furthermore, as mentioned in Sulai et al. (2023); Bloch and Kalia (2024), employing local multi-sensor arrays to measure the curl of the local magnetic field may extend the accessible frequency range to approximately 1 kHz.
Acknowledgement. We thank Saarik Kalia and Daniel Gavilan-Martin for the valuable discussions on the data analysis. L.W. is supported by the NNSFC under Grant No. 12275134 and No. 12335005. This work is supported by the National Key Research and Development Program of China (Grant No. 2022YFA1605000), and Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2024ZD0302300). J.S. is supported by Peking University under startup Grant No. 7101302974 and the National Natural Science Foundation of China (NNSFC) under Grants No. 12025507, No.12450006. J.G. is supported by the NNSFC under Grant No. 12305111. B.Z. is supported by the NNSFC under Grant No. 12275232 and Shangdong Provincial Natural Science Foundation for Distinguished Young Scholars under Grant No. ZR2025QA20. Q.Y. is supported by the Project for Young Scientists in Basic Research of Chinese Academy of Sciences under Grant No. YSBR-061.
@@
Ariel Arza,1,2 Yuanlin Gong,1 Jun Guo,3 Xiaofei Huang,4 Jing Shu,5,6,7 Hongliang Tian,1 Wenyu Wang,8 Kai Wei,4,9 Lei Wu,1,2 Mingming Xia,10 Jin Min Yang,11,12 Qiang Yuan,13,14 Yang Zhang11 Yi Zhang,13,14 Bin Zhu,15
@@
13Key Laboratory of Dark Matter and Space Astronomy, Purple Mountain Observatory, Chinese Academy of Sciences, Nanjing 210008, China